OpenStudy (anonymous):

Evaluate v(w + 3) for v(w) = 2w - 1

OpenStudy (campbell_st):

just replace w in the equation v(w) = 2w -1 with w + 3 so you have v(w +3) = 2(w +3) -1 just distribute and simplify for the answer

OpenStudy (anonymous):

Say, N=w+3, then v(N)=2N-1. substituting back for w you would then get v(w+3)=2(w+3)-1 =2w+6-1=2w+5

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

v is a function not a variable

hero (hero):

lol, I knew that...

OpenStudy (campbell_st):

why, does it change the answer...?

hero (hero):

I can see how one could get confuzzled by it.

OpenStudy (anonymous):

in this case v is known as "a function of w". w can take any value. What they are saying here is that for any w we throw in were gonna add 3 to it. so if w was 1 then it would be 4 and so on. You could just call w+3=x and then say v(x)=2x-1, and if you know how to solve systems of equations then x=w+3 you could solve it that way. OR you can just substitute w+3 for w and distribute both will work

OpenStudy (anonymous):

okay thanks